OpenStudy (anonymous):

What are the vertices of the hyperbola given by the equation y2 – 9x2 + 22y – 72x – 32 = 0? (-3, -11) and (-5, -11) (-1, -11) and (-7, -11) (-4, -8) and (-4, -14) (-4, -10) and (-4, -12)

OpenStudy (anonymous):

had u tried to take it to standar form by completing the square?

OpenStudy (jdoe0001):

$$ y^2-9x^2+22y-72x-32=0\\ (y^2+22y)-(9x^2-72x)-32=0\\ (y^2+22y)-9(x^2-8x)-32=0\\ \text{now completing the perfect square trinomials}\\ (y^2+22y+\boxed{121})-9(x^2-8x+\boxed{16})-32=0 $$ can you take it from there?

OpenStudy (jdoe0001):

$$ y^2-9x^2+22y-72x-32=0\\ (y^2+22y)-(9x^2-72x)-32=0\\ (y^2+22y)-9(x^2-8x)-32=0\\ \text{now completing the perfect square trinomials}\\ (y^2+22y+\boxed{121})-9(x^2-8x+\boxed{16})=32+\boxed{121}-9\times\boxed{16} $$ rather